# Verifying LLM Inference to Detect Model Weight Exfiltration

### v2

@@ 2.3 Non-determinism in Machine Learning
Fixing the random seed (even with temperature T=0T{=}0) does not guarantee identical outputs. Non-determinism arises because floating-point arithmetic is non-associative ($(a{+}b){+}c\neq a{+}(b{+}c)$): different GPU kernels or shapes can accumulate tensors in different orders, yielding different values. Batch variance, where kernel computation strategies depend on the batch size, which depends on the concurrent load, is the most common source of numerical noise. Inference optimizations further shift computation orders—batch size, attention chunking/prefix caching, and cache reuse—and MoE models add routing variability that can depend on load and tiny numerical perturbations [he2025nondeterminism].
@@
Identical open-weight models have shown cross-provider performance discrepancies, consistent with differences in templates, quantization/serving stacks, and the variability sources above [willison_nondeterminism].
Reproducing Non-Reproducibility
To illustrate non-determinism in machine learning inference, we query the OpenAI API with a single fixed prompt, using identical seed and sampling parameters (temperature $0.7$, seed $1234$) across 20,00020{,}000 completions. We tokenize each completion (using whitespace) and construct a token-level prefix tree: each node $v$ represents a prefix $y_{1:t}$, and the fork factor $\phi(v)$ is the number of distinct next-tokens observed. On gpt-4o-mini, we observe that non-determinism is sparse but present: the vast majority of nodes have $\phi(v)=1$ (deterministic), while divergences concentrate at a small number of token positions (Fig. 2).
(a) Token-level prefix tree for 1,0001{,}000 seeded completions (first $100$ positions shown). Red nodes mark fork points (nodes with >1>1 children).
(b) Distribution of fork factors (children per node). Counts are on a log scale.
Figure 2: Seeded non-determinism on a single prompt with OpenAI 4o-mini.
2.4 Steganography and Watermarking
A steganographic scheme is a method for communicating covertly over a channel by embedding hidden information into otherwise ordinary-looking messages. The underlying idea is that the channel already produces messages with some natural distribution (e.g., images, documents, or network traffic). A steganographic encoder uses a random key and generates messages so that they remain consistent with this distribution while carrying additional hidden content decodable only to someone with the key. These messages are often called covertexts.
For instance, in the context of language models, which sample from a probability distribution over tokens, one could encode 1 bit per token by choosing either the most likely or second most likely token based on the hidden bit value. This would still produce natural-looking text, as both tokens will typically have high probability under the model.
Watermarking is a related technique often used to encode a single bit of information within a covertext – $1$ if the watermark is present and 0 if the watermark is not. The embedded information can serve many purposes, such as asserting ownership, tracing provenance, or verifying authenticity. However, importantly, the field of steganography and watermarking can be thought of on a spectrum, where a steganographic message is encoded as a sequence of watermarks. This observation allows us to pull from theoretical bounds in the watermarking literature, as the theory literature for watermarking is generally written with asymptotics, and thus implicitly works in the “huge message regime.”
There is a recent but long line of work in language model watermarking [Aar22, ICML:KGWKMG23, COLT:ChrGunZam24, CIC:FGJMMW23, ARXIV:KTHL23, pseudo_random_error_correcting, golowich2025editdistancerobustwatermarks, SP:CHS25]. However, most works operate in a model that require the watermarker and detector to share only a small key with the goal of making the watermark as robust as possible. This means that even the highest throughput schemes only embed about $1$ bit per “block,” where a block is text that is being watermarked [^2] The length of a “block” depends on the scheme and security parameter, but often requires about 128-bits of empirical entropy. [SP:CHS25]. In contrast, language model steganography [CCS:KJGR21] has a higher throughput rate, since it assumes text is unchanged. These schemes can almost reach capacity, embedding as many bits as the entropy of the response.
In this work, we adopt the language of steganography, with minor modifications listed below. For other cryptographic objects, such as pseudorandom functions, hash functions, etc. we refer the reader to definitions provided in Katz and Lindell [KatLin14].
  1. 1.
    Model inputs (Prompts) $I$: this is what a user will send as a request to the model.
  2. 2.
    Cover-text (model outputs) $O$:, this is what a user would ordinarily receive from correct inference, and this is the channel over which an adversary will attempt to exfiltrate data.
  3. 3.
    Seed $\sigma$: alongside every input output-pair, generation is tied to a seed that is used during the sampling process.
  4. 4.
    External Adversary $\mathcal{A}_{E}$: An adversary who interacts with the model only through its input–output interface, and who has full control over model inputs $I$, and can only observe its outputs.
  5. 5.
    Internal Adversary $\mathcal{A}_{I}$: An adversary with direct access to the model internals (e.g., weights, gradients, or hidden states), who may attempt to exfiltrate information by subtly influencing the model’s outputs, but cannot manipulate model inputs.
  6. 6.
    Warden $\mathcal{W}$: An observer who monitors the channel (model inputs and outputs) and attempts to distinguish between benign inference outputs and outputs that contain hidden.
Definition 2.1 (Steganographic Scheme).
Let $\ell_{\mathrm{msg}},\ell_{\mathrm{cov}}\in\mathbb{N}$ be integers and $\mathcal{T}$ be a set of tokens (a.k.a. alphabet). Let $\mathcal{D}$ be the passive channel distribution with support $X\subseteq\mathcal{T}^{\leq\ell_{\mathrm{cov}}}$. A steganography scheme for $\mathcal{D}$ is a pair of stateful (possibly randomized) algorithms $\mathsf{Steg}=(\mathsf{Enc},\mathsf{Dec})$ with the following syntax:
  • $\mathsf{Enc}(K,\mathsf{st},m)\to c$, takes a key $K\in\{0,1\}^{k}$, a state $\mathsf{st}$, a message $m\in\{0,1\}^{\leq\ell_{\mathrm{msg}}}$ and outputs a covertext $c\in\mathcal{T}^{\leq\ell_{\mathrm{cov}}}$.
  • $\mathsf{Dec}(K,\mathsf{st},c)\to m$, takes a key $K\in\{0,1\}^{k}$, a state $\mathsf{st}$, a covertext $c\in\mathcal{T}^{\leq\ell_{\mathrm{cov}}}$ and outputs a message $m\in\{0,1\}^{\leq\ell_{\mathrm{msg}}}$.
For a steganography scheme to be useful, it must be correct, meaning that for any $\mathsf{st}$ and $m$, $\Pr_{K}[\mathsf{Dec}(K,\mathsf{st},\mathsf{Enc}(K,\mathsf{st},m))=m]=1-\mathrm{negl}(k).$ When, this probability is $1$, we say the scheme is perfectly correct. It must also be secure, meaning that for any $\mathsf{st}$ and $m$, $D_{\mathrm{KL}}(\mathcal{D}_{\mathsf{Steg}}\|\mathcal{D})=\mathrm{negl}(k),$ where $\mathcal{D}_{\mathsf{Steg}}$ is the covertext distribution generated as $K\overset{{\scriptscriptstyle\mathdollar}}{\leftarrow}\{0,1\}^{k}$ and $c\overset{{\scriptscriptstyle\mathdollar}}{\leftarrow}\mathsf{Enc}(K,\mathsf{st},m)$.
The principal measure of efficiency that we are interested in is the rate of a scheme $\mathsf{Steg}$. This is defined as the message length $\ell_{\mathrm{msg}}$ divided by the average covertext length, which we calculate as $\mathbb{E}_{c\sim\mathcal{D}}[|c|]$, since a secure scheme cannot differ significantly from this expectation. Therefore, we write $R=\ell_{\mathrm{msg}}/\mathbb{E}_{c\sim\mathcal{D}}[|c|]$. [^3] This is assuming that one requires correctness for a message distribution which is chosen uniformly at random and is therefore incompressible.
Note that there are many other variants of steganography, including stateless, public-key, robust, etc. [JOC:AP02, C:HLvA02, IDK:vAH04]. However, in our context, we wish to give a potential exfiltrator as much theoretical power as possible. So we assume that both endpoints are stateful and can have an arbitrarily long secret key in common.
2.4.1 Relevant Results in Steganography
Theorem 2.2 (Steganographic limitations).
Let $\mathsf{Steg}$ be a steganography scheme for $\mathcal{D}$ and $\mathbf{M}$ be a random variable with $H(\mathbf{M})=\ell$. Let $\mathcal{D}_{\mathsf{Steg}}$ be the covertext distribution generated as $K\overset{{\scriptscriptstyle\mathdollar}}{\leftarrow}\{0,1\}^{k}$ and $c\overset{{\scriptscriptstyle\mathdollar}}{\leftarrow}\mathsf{Enc}(K,\mathsf{st},\mathbf{M})$. Then, $\mathsf{Steg}$ cannot be perfectly correct over the choice of $\mathbf{M}$ if $\ell>\lceil H(\mathcal{D}_{\mathsf{Steg}})\rceil.$
Proof.
This theorem is an immediate consequence of Shannon’s source coding theorem [shannon1948mathematical, shannon1959coding]. This theorem states: (i) that no random variable $X$ can be encoded with fewer bits than $H(X)$ in expectation and (ii) that there exists an encoding of $X$ with expected length at most $\lceil H(X)\rceil$. Therefore, if $\lceil H(\mathcal{D}_{\mathsf{Steg}})\rceil<H(\mathbf{M})$ and $\mathsf{Steg}$ is correct with probability $1$, then one could encode $\mathbf{M}$ as the covertext, using a random key $K$ and random state $\mathsf{st}$, $\mathbf{C}\leftarrow\mathsf{Enc}(K,\mathsf{st},\mathbf{M})$ and decoded with $\mathsf{Dec}(K,\mathsf{st},\mathbf{C})$. Then, based on the source coding theorem, $\mathbf{C}\sim\mathcal{D}_{\mathsf{Steg}}$ (and therefore $\mathbf{M}$) can be encoded with $\lceil H(\mathcal{D}_{\mathsf{Steg}})\rceil<H(\mathbf{M})$ bits, which is a direct contradiction. ∎
In contrast to the above theorem and to illustrate the difficulty of preventing covert exfiltration, we recall a well-known theorem from [C:HLvA02] stated in Theorem 2.3. This shows that assuming classical cryptography exists at all, i.e., one-way functions exist, then one can construct a steganographic scheme that is near-optimal.
Theorem 2.3 (Due to [C:HLvA02]).
Let $\mathcal{P}$ be an efficiently-computable policy with a false positive rate $\alpha$ over $\mathcal{D}$. Then, assuming a $\delta$-secure pseudorandom function $\mathsf{PRF}$ exists, there is an inefficient steganography scheme $\mathsf{Steg}$ for $\mathcal{D}$ which, on every message that is $H_{\mathrm{min}}(\mathcal{D})-O(\log(1/\varepsilon))$ bits long, is correct with probability at least $1-\varepsilon$ and is caught with probability at most $\alpha+\delta$ by $\mathcal{P}$.
Finally, we state a known result from information theory in Theorem 2.4. This result gives a lower bound on the number of messages that must be sent through independent channels, each of capacity $C$. This is a useful result to recall for the bounds we derive later in Section 4.6.
Theorem 2.4 (Capacity limitations).
Let $X$ be a random variable with support of size $q$ (and entropy $H(X)=\log q$) and $Y^{n}$ be $n$ messages sent through a channel of capacity $C$. Then, in order to transmit $X$ with error at most $\varepsilon$, requires that $n\geq\big(\log q-h(\varepsilon)-\varepsilon\log(q-1)\big)/C.$ Notably, when $Y$ has alphabet size $q^{\prime}$ and is uniformly random with probability $p$, then we obtain
Proof.
$H(X\mid Y^{n})\leq h(\Pr[E])+\Pr[E]\cdot\log(q-1).$
Using the mutual information between $X$ and $Y^{n}$,
2.5 Model Weight Exfiltration
In the context of discussing model-weight exfiltration, a recent framework [nevo2024securing] introduces the notions of security levels from 1 to 5, which increase in defensive capacity. SL1 is a system that can likely thwart amateur attempts. SL2 is a system that can likely thwart most professional opportunistic efforts by attackers that execute moderate-effort or nontargeted attacks SL3 is a system that can likely thwart cybercrime syndicates or insider threats. SL4 is system that can likely thwart most standard operations by leading cyber-capable institutions. SL5 is a system that could plausibly be claimed to thwart most top-priority operations by the top cyber-capable institutions.
To set the grounding for this work, we will focus primarily on creating a scheme for moving from SL3 to SL4 and buulding the basis for SL5 security, by moving time to covertly exfiltrate from minutes and hours to days and years.
@@ 3 Formulating the Model Weight Exfiltration Security Game
In this section, we provide a concrete specification for the exfiltration problem that we solve in this paper. At a high-level, we consider the setting of LLM inference where there is an inference server and a chat server; we assume the inference server is fully corrupted by an adversary whose goal is to exfiltrate the inference model weights to an outside party, with whom they can coordinate. We assume that the adversary does not have access to any side-channels, eliminating exfiltration techniques such as using timing channel attacks, uploading the weights to a server through a developer’s laptop, or physically stealing the device. In many ways, this task is not different from a traditional security problem attempting to prevent any kind of exfiltration; there are several main differentiating factors associated with model-weight exfiltration.
@@
There are two main advantages to the defender: (i) Model weights are large; (ii) The type of operations and data that a model handles and outputs is largely predictable (e.g. matrix multiplications and text data).
@@ 3.1 Threat Model: Baseline Trust Assumptions
We consider an idealized setting of the model-weight exfiltration problem. In particular, we do not attempt to defend against adversaries with unrestricted physical access to hardware (e.g., physically stealing servers or storage devices to extract model weights). Our focus is on protecting model parameters in environments where security is enforced through software, networking, and operational controls. A more detailed discussion of these assumptions is provided in Appendix D.
@@
3.2 Security Game
Next, we outline a more formal security game to identify the problem that we are trying to solve. The formal description is given in Experiment 3.2, which is parameterized primarily by four integers variables:
  • $n$, the number of bits which the adversary is trying to exfiltrate,
  • $M$, the number of prompt/response pairs the adversary has control over,
  • $B_{I}$, the number of inspections the warden may issue across the $M$ messages,
  • $B_{A}$, the number of alarms that the warden may raise across the $M$ messages.
At a high level, the adversary’s goal in the game is to coordinate between its internal algorithm $\mathcal{A}_{I}$, which has access to the challenge message, and its external algorithm $\mathcal{A}_{E}$, which is tasked with learning the challenge message. At each of the $M$ steps, the internal adversary may choose to either generate an honest response from a latent distribution or a dishonest response from its chosen distribution. The wardens wins only if it raises the alarm on these dishonest messages or if $\mathcal{A}$ fails to learn the challenge message.
{experiment}
[Exfiltration-with-Auditing Experiment] Parameters. Let $\theta$ denote the deployed model weights. Fix corpus size $|\mathcal{M}|$ (total input messages) and outputs $O$, where $O$ is the valid outputs of model $\theta$ on inputs $m\in\mathcal{M}$. $O\leftarrow\theta(m)$. Let $B_{I}\in\mathbb{N}$ (inspection budget), $B_{A}\in\mathbb{N}$ (alarm budget). Let $\mathbf{m}$ be a hidden uncompressible message of length $n:=|\mathbf{m}|$ bits (modeled as a random bit-string) that $\mathcal{A}$ attempts to exfiltrate (by default one can imagine this being the model weights themselves). Adversary $\mathcal{A}=(\mathcal{A}_{I},\mathcal{A}_{E})$. Warden $\mathcal{W}$ may inspect up to $B_{I}$ outputs (inspection budget) and raise up to $B_{A}$ alarms (auditing budget). All post-start communication is restricted to the model I/O channel.
$\mathsf{ExfilGame}_{\mathcal{A},\mathcal{W}}(n,M,B_{I},B_{A}):$
  1. 1.
    Setup. Compute starting states $\mathsf{st}_{I}$ and $\mathsf{st}_{E}$ for $\mathcal{A}_{I}$ and $\mathcal{A}_{E}$, which can contain shared randomness or other shared strategy information. Initialize $q_{I}=q_{A}=0$. Sample random $n$-bit message $\mathbf{m}$.
  2. 2.
    Interaction. For $t=1,\dots,M$:
    1. (a)
      $\mathcal{A}_{E}(\mathsf{st}_{E})$ submits query $(x_{t},\sigma_{t})$, where $\sigma_{t}$ is the random seed for response generation.
    2. (b)
      $\mathcal{A}_{I}(\mathbf{m},\mathsf{st}_{I},x_{t},\sigma_{t})$ either outputs an (honest) response $y_{t}$ or chooses a (dishonest) response $y_{t}$, and updates $\mathsf{st}_{I}$.
    3. (c)
      $\mathcal{W}$ may inspect $(x_{t},y_{t},\sigma_{t})$, incrementing $q_{I}$. Then, $\mathcal{W}(x_{t},y_{t},\sigma_{t})$ may raise an alarm, incrementing $q_{A}$. If the alarm is raised, $y_{t}$ is dishonest, $q_{I}\leq B_{I}$, and $q_{A}\leq B_{A}$ and then $\mathcal{W}$ wins (output 0).
  3. 3.
    Termination. If $q_{I}>B_{I}$ or $q_{A}>B_{A}$, then $\mathcal{A}$ wins (output $1$). Then, let $\mathcal{A}_{E}$ output $m^{\prime}$. If $m^{\prime}=\mathbf{m}$, then $\mathcal{A}$ wins (output $1$). Otherwise, $\mathcal{W}$ wins (output 0).
Definition 3.1 ($(P,n,M,B_{I},B_{A})$-exfiltration secure).
A warden $\mathcal{W}$ is $(P,n,M,B_{I},B_{A})$-exfiltration secure such that for all PPT adversaries $\mathcal{A}=(\mathcal{A}_{I},\mathcal{A}_{E})$,
$\Pr\big[\mathsf{ExfilGame}_{\mathcal{A},\mathcal{W}}(n,M,B_{I},B_{A})=1\big]<P.$
Notice that the adversary in this game could concentrate all of their exfiltration on just one message, hoping that the warden doesn’t inspect it. While this isn’t feasible in practice, it demonstrates that formally, the adversary can always win in the abstract game with probability at least $(M-B_{I})/M$. This serves as the “best” we can hope for in the model as its proposed.
In the next sections, we show that using a carefully designed warden which can have both few false positive flags on honest traffic and rate that exfiltrators cannot exceed with high probability without being caught. We also complement this theoretical analysis with empirical data and modeling that supports our conclusions. This lets us essentially meet the above baseline probability under certain trust assumptions.
We note that the precise form of the security game is a design choice. A higher-level objective in constructing a detection scheme is to minimize the trust assumptions required by the game, while still ensuring that the expected time to successful exfiltration remains above an acceptable threshold.
We expand upon these assumptions in the context of our specific scheme in Section 5. We discuss possible policy-responses to exfiltration-detection in the Appendix F.

### v3

@@ 3.3 Non-determinism in Machine Learning
Despite the amount of reproducibility of the methods above, fixing the random seed (even with temperature T=0T{=}0) does not guarantee identical outputs. Non-determinism arises because floating-point arithmetic is non-associative ($(a{+}b){+}c\neq a{+}(b{+}c)$): different GPU kernels or shapes can accumulate tensors in different orders, yielding different values. Batch variance, where kernel computation strategies depend on the batch size, which depends on the concurrent load, is the most common source of numerical noise. Inference optimizations further shift computation order; batch size, attention chunking/prefix caching, and cache reuse, and MoE models all add routing variability that can depend on load and tiny numerical perturbations [HL25].
@@
Identical open-weight models have shown cross-provider performance discrepancies, consistent with differences in templates, quantization/serving stacks, and the variability sources above [WIL25].
We empirically characterize this non-determinism by querying the OpenAI API with fixed seeds across thousands of completions; the results confirm that divergences are sparse but present, concentrating at a small number of token positions (see Appendix J).
While Inverse-Probability Transform and Gumbel-Max Trick are mathematically deterministic, in practice the logit distribution $\vec{p}_{\theta}(x)$ is produced by a non-deterministic computation. Even with a fixed seed $\sigma$, floating-point imprecision, asynchronous GPU execution, and nondeterministic kernel scheduling can cause small but measurable differences in $\vec{p}_{\theta}(x)$ across runs. This effect arises even in benign settings, independent of any adversarial manipulation, and has been well-documented in prior work [YLD+25]. We refer to this inherent variability as valid non-determinism.
$\vec{p}_{\theta}(x)\sim\mathcal{P}_{\theta}(x),$
where $\mathcal{P}_{\theta}(x)$ represents the stochastic process generating the token-level probabilities. This perspective allows us to define heuristic distance measures between two runs of the same model (with identical seed $\sigma$) by comparing samples from $\mathcal{P}_{\theta}(x)$, illustrated conceptually in Figure 2.
Fixed-Seed Posterior Distribution.
$t=\mathrm{Sample}(p,\sigma).$
This represents the probability, over draws of the underlying probability distribution $p$, that a fixed random seed $\sigma$ yields token $t$. Intuitively, it characterizes how much stochastic variability remains in the sampling outcome even when the seed is fixed, due to non-determinism in the generation of $p$.
$P_{i}(t)=\Pr[T=t\mid X=x_{i}].$
$S_{i}(\sigma)=\min\{\,t\in\mathcal{V}:F_{i}(t)\geq R(\sigma)\,\},$
Intuitively, this distribution captures how much apparent randomness remains even when the seed is fixed. Figures 2(a) and 2(b) visualize the resulting fixed-seed posterior distributions for the inverse probability transform and the Gumbel-Max mechanism, respectively. These distributions form the basis for reasoning about an honest divergence during verification (as opposed to potentially malicious nondeterminism).
Understanding the limits of valid non-determinism is central to limiting the exfiltration rate of adversaries. Under very minimal trust assumptions (Appendix E), the only communication channel open to the adversary is by hiding communication within this honest variability.
@@ 4 Formulating the Model Weight Exfiltration Security Game
In this section, we provide a concrete specification for the exfiltration problem that we solve in this paper. At a high-level, we consider the setting of LLM inference where there is an inference server and a chat server; we assume the inference server is fully corrupted by an adversary whose goal is to exfiltrate the inference model weights to an outside party, with whom they can coordinate. We assume that the adversary does not have access to any side-channels, eliminating exfiltration techniques such as using timing channel attacks, uploading the weights to a server through a developer’s laptop, or physically stealing the device holding the model. In many ways, preventing model theft is not different from a traditional security problem attempting to prevent any kind of exfiltration; there are several main differentiating factors associated with model-weight exfiltration.
@@
There are two main advantages to the defender: (i) Model weights are large; (ii) The type of operations and data that a model handles and outputs is largely predictable, namely: matrix multiplications and text data.
@@
4.1 Security Game
We define a simple extraction game, which we later be used to bound the success probability of an adversary extracting information, in Section 5.3. The formal description is given in Experiment 3, which is parameterized by:
  • $\mathcal{T}$, the token set (i.e., vocabulary or alphabet) of the adversary’s message.
  • $\mathsf{H}$, is a hash or pseudorandom function, modeled as a random oracle for simplicity.
  • $\mathcal{A}_{0}$, $\mathcal{A}_{1}$, $\mathcal{A}_{2}$, the adversary’s preprocessing, encoding, and decoding algorithms respectively.
  • $\mathcal{D}_{\theta,\mathcal{H}}(\mathsf{seed})$, the honest response distribution , determined by the model weights $\theta$, context (or prompt) $\mathcal{H}$, and seed $\mathsf{seed}$. In our context, even after all these are given, the output is a randomized because of “invalid nondeterminism” (Section 3.3). In our security games, we suppress the $\theta$ and $\mathcal{H}$ parameters and treat them as fixed for the given adversary.
  • $\mathcal{W}$ (warden), the policy algorithm to determine whether to flag a message.
  • $n$, the number of bits which the adversary is challenged to exfiltrate (in our setting, this is the length of the model weights).
  • $\ell$, the length of the adversary’s message.
  • $\lambda$, the seed length of the model’s randomness and the security parameter of the system.
Success in this security game is defined by the adversary successfully reconstructing the challenge bits from the message (and any other precomputed material) itself. We define security for computationally bounded adversaries and policies as well as unbounded algorithms.
In our games, we suppress the notation of the model weights $\theta$ and the prompt $P$. We abstract the prompt out of the security game because a practical adversary will have already picked some prompt $P$ on which they will try and perform the exfiltration. And we care about limiting the exfiltration rate on this particular prompt. We also choose to model an adversary trying to exfiltrate some information about the incompressible model weights $\theta$ as the adversary sending and recovering a random bit-string. This allows us to further abstract away the model weights from the security game.
Note this security has been explored in the context of steganography, which we draw on directly. However, we are not interested in proving security against all possible policies/wardens. One can view our results for specific lower bounds for steganography in a particular realistic context.
Figure 3: Outline of the security game described in Experiment 3.
{experiment}
[Exfiltration Experiment] Parameters. Let positive integers $n$ be the challenge bit length, $\ell$ be the maximum response length, and $\lambda$ be the security parameter. Let $\mathcal{T}$ be the token set (vocabulary) of messages and $\mathcal{D}^{\mathsf{H}}(\sigma)$ be a randomized function from $\{0,1\}^{\lambda}$ to $\mathcal{T}^{\leq\ell}$. Then, for a policy $\mathcal{W}$ and adversary $\mathcal{A}=(\mathcal{A}_{0},\mathcal{A}_{1},\mathcal{A}_{2})$, we define the games:
$\mathbf{Exfil}_{\mathcal{W},\mathcal{A},\mathcal{D}}(\lambda,n,\ell)$$\mathbf{Exfil}^{\mathbf{r}}_{\mathcal{W},\mathcal{A},\mathcal{D}}(\lambda,n,\ell)$
  1. 1.
    Initialize a random oracle $\mathsf{H}:\{0,1\}^{*}\to\{0,1\}$
  2. 2.
    $\mathsf{st}_{1},\mathsf{st}_{2}\leftarrow\mathcal{A}_{0}^{\mathsf{H}}(1^{\lambda},1^{n},1^{\ell})$    // preprocessing
  3. 3.
    $X\leftarrow\{0,1\}^{n}$ ; $\sigma\leftarrow\{0,1\}^{\lambda}$   // sample challenge and seed
  4. 4.
    $\hat{T}\leftarrow\mathcal{A}_{1}^{\mathsf{H}}(\mathsf{st}_{1},X,\sigma)$      // generate cover response
  5. 5.
    If $\mathcal{W}^{\mathsf{H}}(\sigma,\hat{T})=1$, then output $\bot$
  6. 6.
[…]
  7. 7.
    If $\hat{X}=X$, then output 11
  8. 8.
    Else, output 0.
These definitions are presented in the random oracle model [BR93], which enables some of our later results. Note that all parties know and have access to $\mathcal{D}^{\mathsf{H}}.$
The first game $\mathbf{Exfil}$ is the standard game, where both parties are aware of the seed that is used to sample the distribution. However, in cases when the seed $\sigma$ is sampled securely (and $\mathcal{A}_{2}$ does not get it directly), we are able to obtain stronger bounds in the more restricted game $\mathbf{Exfil}^{\mathbf{r}}$.
Definition 4.1 ($(\alpha,\varepsilon)$-exfiltration secure).
A policy $\mathcal{W}$ is (computationally) $(\alpha,\varepsilon)$-exfiltration secure if:
  1. 1.
    For all $\sigma$, $\Pr_{T\sim\mathcal{D}^{\mathsf{H}}(\sigma)}\big[\mathcal{W}^{\mathsf{H}}(\mathsf{seed},T)=1\big]\leq\alpha,$ i.e., $\mathcal{W}$ has false positive rate $\alpha$,
  2. 2.
    and for all (PPT) adversaries $\mathcal{A}$, $\Pr\big[\mathbf{Exfil}_{\mathcal{W},\mathcal{A},\mathcal{D}}(\lambda,n,\ell)=1\big]-2^{-n}\leq\varepsilon.$
Note that $\alpha$ and $\varepsilon$ are, in general, functions of the parameters $\lambda$, $n$, $\ell$, or even $\mathcal{D}$. If the above holds only for $\mathbf{Exfil}^{\mathbf{r}}$ rather than $\mathbf{Exfil}$, then we say it $\mathcal{W}$ is secure against seed agnostic decoders.
Where relevant, we will also consider adversaries with restricted access to the random oracle $\mathsf{H}$.
@@ 4.2 Threat Model: Baseline Trust Assumptions
We consider an idealized setting of the model-weight exfiltration problem. In particular, we do not attempt to defend against adversaries with unrestricted physical access to hardware (e.g., physically stealing servers or storage devices to extract model weights). Our focus is on protecting model parameters in environments where security is enforced through software, networking, and operational controls. A more detailed discussion of these assumptions is provided in Appendix E.